OpenStudy (anonymous):

If 6m + 4 = 8m, then 4m

OpenStudy (texaschic101):

6m + 4 = 8m -- subtract 6m from both sides 4 = 8m - 6m 4 = 2m -- divide both sides by 2 4/2 = m 2 = m then 4m = 4(2) = 8

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

6m+4=8AM 4=8-6 4=2 4*2 8